dx/y+1 =dy/X+1=dz/z


1
Expert's answer
2021-06-16T13:19:23-0400

"\int(x+1)dx=\int(y+1)dy"


"x^2\/2+x=y^2\/2+y+c_1"


"\frac{dx-dy}{y-x}=\frac{dz}{z}"


"-ln(x-y)=lnz+lnc_2"


"\frac{1}{x-y}=c_2z"


"F(x^2/2+x-y^2/2-y,\frac{1}{z(x-y)})=0"